# Rydberg superatoms: An artificial quantum system for quantum information processing and quantum optics

### v1

@@ IV.1 Quantum logic gate
For multiatom ensemble qubits, Saffman’s group demonstrates coherence control between two ensemble qubits and achieves Rydberg blockade between them with a fidelity of 0.89(1).Ebert et al. (2015) However, fluctuations in the collective Rabi frequency could greatly affect the ensemble qubits operation fidelity. Thus, a scheme is proposed to simulate quantum process tomography to characterize Rydberg gates based on collective states using optimal stimulated Raman adiabatic passage pulses. Beterov et al. (2016b) Recently, significant progress has been made in achieving entanglement gates through coherent manipulation of deterministically assembled arrays, employing local optical control to achieve entanglement using interatomic interactions, although individual addressing of atoms remains a challenge experimentally. Zeng et al. (2017); Scholl et al. (2023); Evered et al. (2023); Ma et al. (2023) Additionally, due to the excellent transmission characteristics of photons but the lack of strong interactions, entangling gate schemes for photon-atom hybridization have been proposed. Garcia-Escartin and Chamorro-Posada (2012); Paredes-Barato and Adams (2014); Hao et al. (2015); Das et al. (2016); Sárkány, Fortágh, and Petrosyan (2015) Specifically, the combination of a Rydberg blockade and optical cavities effectively enhances the gate robustness, Das et al. (2016) while a two-photon interference through virtual photon exchanges renders the internal state dynamics of two Rydberg atomic qubits insensitive to the thermal photon state. Sárkány, Fortágh, and Petrosyan (2015)
@@
Figure 12: Level scheme for EIT storage of the control qubit (a) and target qubit (b). (c) The schematic of the experimental setup. Reprinted with permission from Tiarks et al., Nature Physics 15(2), 124-126 (2019). Copyright 2019 Springer Nature.Tiarks et al. (2019)
@@
Tiarks et al. achieved the first experimental realization of a photon-photon quantum gate using Rydberg interactions, as documented in their study.Tiarks et al. (2019) By harnessing both EIT and Rydberg interactions within an ultracold atomic ensemble, they effectively encoded quantum information carried by photons into specific atomic states, as depicted in Fig. 12(a). This mapping facilitates the control of photon loss through post-selection, thereby mitigating its detrimental impact on quantum information and ensuring resilience against errors stemming from photon loss. Notably, the photon interaction occurs exclusively when they share the same polarization $|L\rangle$ due to the Rydberg component, and their interaction during temporal and spatial overlap results in the accumulation of a conditional phase shift, thus realizing a controlled phase flip gate, as illustrated in Fig. 12(b). By altering the basis, this gate can be converted into a controlled-NOT (CNOT) gate. The fidelity of the controlled-NOT truth table and the entangling-gate fidelity, post-selected upon the detection of control and target photons, are reported to be 70(8)and 63.7(4.5)%, respectively. This technique employs effectively Rydberg interactions and EIT, allowing for strong, long-range interactions between photons while also solving the difficulty of restricted photon interactions. By mapping Rydberg interactions onto photons, this gate approach stores and manipulates quantum information, demonstrating high fidelity and potential advances in quantum communication systems and quantum networking applications.
@@
Figure 13: Schematic diagram of converting a circuit into an atomic arrangement, which consists of two types of atoms, and superatoms are used as impurities. Reprinted with permission from Cesa et al., Phys. Rev. Lett. 131(17), 170601 (2023). Copyright 2023 American Physical Society.Cesa and Pichler (2023)
@@ IV.2 Quantum entanglement
Quantum entanglement plays a crucial role in quantum computing and quantum information processing. Early experiments of entangled states have focused on photons and ions. Prevedel et al. (2009); Wieczorek et al. (2009); Häffner et al. (2005) However, these schemes face challenges in scaling to multiqubit entanglement and have a low success probability. Rydberg ensembles offer a promising solution to overcome these limitations and realize scalable entanglement of neutral atoms. Atom vapors serve not only as quantum memory, but also as “good” qubits due to strong Rydberg interactions. High-efficiency cluster states have been created via single-qubit measurements with Rydberg ensembles. Zwierz and Kok (2009) Utilizing a fast entanglement generation scheme with a Lipkin-Meshkov-Glick-type Hamiltonian, a 700-atom cat state has been generated with a 50-fold improvement in the size of the entangled state. Khazali (2018) To further enhance fidelity, Gärttner proposes a scheme to herald the entangled state by detecting an ion, achieving near-unity fidelity. Gärttner (2015) Additionally, addressing decoherence concerns, one can combine Zeno dynamics and shortcuts to adiabaticity with Rydberg superatoms, encoding quantum information in the collective state of superatoms to rapidly realize entangled states, including Bell states and $W$ states. Ji et al. (2020) Similarly, fast conversion of three-particle Dicke states to four-particle Dicke states with Rydberg superatoms can be implemented based on Zeno dynamics and shortcuts to adiabaticity, offering high fidelity and robustness against decoherence. Liu et al. (2023b) In Rydberg ensembles, large-scale entangled states can also be prepared by sending two Rydberg superatoms belonging to two different atomic $W$ states into a cavity without qubit loss. Zhang et al. (2022a)
@@
Figure 14: Entanglement between Rydberg superatom and photon. (a) Schematic diagram of the entanglement of light with optical atomic excitation. Mapping the collective atomic excitation into quantum fields $|\phi\rangle_{A,B}$ by implementing resonance on the $|r\rangle$ and $|e\rangle$ used laser fields $\Omega_{A,B}$. Reprinted with permission from Li et al., Nature 498(7455), 466-469 (2013). Copyright 2013 Springer Nature.Li, Dudin, and Kuzmich (2013) (b) Experimental preparation of semideterministic atom-photon entanglement. Reprinted with permission from Li et al., Phys. Rev. Lett. 123(14), 140504 (2019). Copyright 2019 American Physical Society.Li et al. (2019b)
@@
Photons are excellent carriers for quantum information, and the entanglement between atom and photon holds promise for achieving long-range quantum information distribution. Most ensemble-based atom-photon entanglement schemes are probabilistic and suffer from limited scalability. In 2013, Li et al. introduced a new protocol to generate entanglement between light and optical atomic excitations with intrinsic determinism. Li, Dudin, and Kuzmich (2013) In this scheme, the dephasing of optical atomic coherence is suppressed by state-insensitive confinement of ground and Rydberg states of ultra-cold atomic gas trapped in a one-dimensional optical lattice at 1004 nm. This confinement eliminates the differential energy shift between the ground state and the Rydberg state $\delta U=U_{r}(r)-U_{g}(r)$ and maintains coherence between them. Such a scheme has significant implications for extending the ground-Rydberg coherence and building functional Rydberg quantum networks. Fig. 14(a) illustrates the three-step process for the entanglement protocol: (i) Excitation of the Rydberg atomic ensembles from the ground state $|G\rangle$ to the single excited state $|R\rangle$. (ii) Generation of the entanglement state $|R\rangle\rightarrow|R\rangle|0\rangle_{A}+|G\rangle|1\rangle_{A}$ through a retrieval field $\Omega_{A}$. (iii) Verification of the atom-photon entanglement by analyzing the correlations of photoelectric detection events. This scheme is compatible with quantum logic operations and long-term storage of quantum information, paving the way for quantum networks that enable secure communication and distributed quantum computing.
@@
In 2019, an alternative atom-light entanglement protocol was experimentally demonstrated, leveraging Rydberg blockade and Raman coupling techniques. Li et al. (2019b) The manipulation of atoms within a confined volume of approximately 10 $\mu$m in all three dimensions ensures high effectiveness of Rydberg blockade, with an atomic density of approximately $10^{11}/\rm cm^{3}$. In the schematic depicted in Fig. 14(b), the atomic ensemble is initially prepared in a ground state $|g\rangle$, following which Rydberg blockade is employed to sequentially generate two collective excitations—one in a ground state $|s\rangle$ and the other in a Rydberg state $|r\rangle$. These collective excitation states, referred to as two superatoms, possess distinct momenta, are then entangled by incorporating the momentum degree of freedom and interfering them through Raman coupling. Finally, one excitation is retrieved using a read beam, converting the momentum-entangled atomic state into entanglement between the polarization of the single photon and the momentum of the remaining atomic excitation, thereby achieving atom-photon entanglement. The high fidelity measurement of entanglement between the polarization of the single photon and the momentum of the atomic excitation underscores the practicality of employing this approach for quantum information processing tasks like quantum repeaters and networks. Although this scheme is semi-deterministic compared to the approach in Ref. Li, Dudin, and Kuzmich, 2013, it remains immune to phase fluctuations and is devoid of high-order or vacuum components, which are essential for remote entanglement generation. After that, the same group presented a report detailing the deterministic creation of entanglement between an atomic ensemble and a single photon.Sun et al. (2022) This achievement was accomplished through a cyclical process termed “retrieving and patching”, which was applied sequentially to the Rydberg levels. By leveraging the Rydberg blockade effect, they ensured that a full excitation was generated only if a preceding excitation existed in one of the Rydberg levels. This process led to the creation of a photon in a delayed temporal mode, entangled with the Rydberg levels hosting the collective excitation. Verification of the entanglement between the atomic ensemble and the single photon was conducted by retrieving the atomic excitation as a second photon and performing correlation measurements. The results yielded an entanglement fidelity of 87.8%.
@@
The entanglement between photons holds tremendous potential across various fields, such as metrology and quantum computing. However, the direct interaction between photons is inherently weak, necessitating certain processing steps to induce their interaction. Heisenberg and Euler (1936); Busche et al. (2017); Firstenberg et al. (2013) Numerous schemes have been proposed for the preparation of multiphoton entanglement, including spontaneous parametric downconversion Pan et al. (2012); Zhong et al. (2018) and the entangling of single photons through atom-photon interactions. Gheri et al. (1998); Schön et al. (2005) However, these schemes often suffer from drawbacks such as poor scalability of the photon number, complex setups, and other requirements on the system. Recently, Yang et al. proposed a novel scheme to generate multiphoton entanglement with a Rydberg superatom. Yang et al. (2022a) In their approach, an atomic ensemble is used, where an atomic entanglement state $|\Psi\rangle_{\mathrm{a}}$ is prepared by applying collective pulses with the Rydberg blockade mechanism. Subsequently, a “retrieving and patching” sequence is applied to $|r_{1}\rangle$, causing the Rydberg superatom to emit a photon and recreate the excitation. This sequence is repeated for $|r_{2}\rangle$, resulting in the creation of the entanglement state $(|0\rangle_{1}|1\rangle_{2}|L\rangle+|1\rangle_{1}|0\rangle_{2}|R\rangle)/\sqrt{2}$. By iterating these processes $m-1$ times and retrieving the two atomic qubits, the GHZ-type multiphoton entanglement $(|L\rangle^{\otimes m}+|R\rangle^{\otimes m})/\sqrt{2}$ can be obtained, where $|L\rangle$ and $|R\rangle$ represent a photon in the late and early modes, respectively. To enhance the performance of multiphoton entanglement, the scheme employs a low-finesse cavity to improve photon collective efficiency and polarizing beamsplitters with a Pockels cell in measurement to minimize exposure. The fidelity of the final prepared three- and six-photon GHZ states is $82.9\pm 0.3\%$ and $61.8\pm 2.6\%$, respectively.
@@
Figure 15: Photonic entanglement filter with Rydberg superatoms. (a) Illustration of Rydberg entanglement filter protocol. (b) Extracting high-fidelity entangled state from input state with low fidelity. Reprinted with permission from Ye et al., Nature Photonics 17(6), 538–543 (2023). Copyright 2023 Springer Nature.Ye et al. (2023)
@@ IV.3 Quantum simulation
Quantum simulation serves as a powerful tool for understanding and predicting many-body dynamics, allowing access to a wide range of phenomena, including thermal equilibrium relaxation in closed systems and non-thermal behavior in integrable and disordered systems. A particularly promising platform for investigating the dynamics of strongly interacting many-body quantum systems is provided by coherent laser-driven Rydberg atomic gases.Weimer (2015); Lechner and Zoller (2015); Ding et al. Currently, spin systems based on Rydberg atoms have emerged as a promising platform for quantum simulation because of their superior interaction flexibility and strength, high controllability, and effective isolation from the environment.
One strategy suggested by Sun and Robicheaux involves simulating Rydberg gases and exploring two-body correlation by driving the superatoms, which can be subdivided into smaller regions called pseudoatoms, with a laser having a Rabi frequency following a Poissonian distribution. Sun and Robicheaux (2008) This method incorporates the classical correlation and provides a more accurate depiction of the Rydberg gas compared to the mean field approximation. Using the superatom picture, Olmos and colleagues conducted extensive research on fermionic collective excitations and thermalization in a symmetric quasi-one-dimensional ring lattice gas of Rydberg atoms. Olmos, González-Férez, and Lesanovsky (2009); Olmos, Müller, and Lesanovsky (2010) In the study of Refs. Pohl, Demler, and Lukin, 2010; van Bijnen et al., 2011, the Rydberg superatoms serve as building blocks for the creation of Rydberg crystals using tailored laser excitation schemes. These crystals are used as representations of dilute metallic solids with tunable lattice parameters, allowing access to a wide range of fundamental phenomena in condensed matter physics.
Figure 16: Average density of holes $\rho_{h}$, Mandel $Q$ parameter for the total number of holes, and amplitudes of the density–density correlation function $g^{\left(2\right)}\left(d\right)$ for the periods of $d=2a$ (blue circles) and $d=3a$ (red circles), versus the hopping rate $\Gamma_{t}$. Reprinted with permission from Letscher et al., New J. Phys. 19(11), 113014 (2017). Copyright 2017 Institute of Physics, licensed under a Creative Commons Attribution 3.0 licence.Letscher, Petrosyan, and Fleischhauer (2017)
@@
Figure 17: Ground state population and corresponding Fourier transform in $N$-body ($N=3,4$) configuration. Reprinted with permission from Kim et al., PRX Quantum 1(2), 020323 (2020). Copyright 2020 American Physical Society, licensed under a Creative Commons Attribution 4.0 International license.Kim et al. (2020)
@@
Figure 18: The numerically calculated probabilities $q_{Ry}\left(n\right)$ to excite $n$ Rydberg atoms in the mesoscopic atomic ensemble with $\bar{N}=7$ atoms, randomly distributed in the optical dipole trap with different radius $r$. Reprinted with permission from Beterov et al., Phys. Rev. A 90(4), 043413 (2014). Copyright 2014 American Physical Society.Beterov et al. (2014)
@@
One of the challenges in observing the Rydberg blockade is to obtain a precise measurement of the quantity of Rydberg atoms. Due to low detection efficiencies, accurately determining the number of excited Rydberg atoms becomes problematic, which is essential for exploiting the Rydberg blockade in quantum information applications. Beterov et al. investigated the complex dynamics of Rydberg excitations in atomic ensembles influenced by a classical electromagnetic field.Beterov et al. (2014) They demonstrated the observation of collapses and revivals in the average count of Rydberg excitations, resembling the behavior predicted by the Jaynes-Cummings model. Through the analysis of randomly loaded optical dipole traps, they predicted the occurrence of collapses and revivals in Rabi oscillations, emphasizing how the oscillation frequency depends on the number of interacting atoms. Moreover, the study examines the impact of finite interaction strengths and laser linewidth on the detectability of the revivals. Considering a randomly loaded optical dipole trap with an average atom number of $\bar{N}=7$, data in Fig. 18 illustrate the excitation dynamics in a Rydberg state in the ensemble. In Fig. 18(a), the simulated dynamics of the probability of single-atom Rydberg excitation $q_{Ry}\left(1\right)$ for $r=2~{}\mu\rm m$ show collapses and revivals, while $q_{Ry}\left(2\right)$ approaches zero, indicating an effective Rydberg blockade. However, with increasing radius of the optical dipole trap, the revivals in $q_{Ry}\left(1\right)$ are diminished, and the probability of exciting two Rydberg atoms $q_{Ry}\left(2\right)$ increases, suggesting the breakdown of the Rydberg blockade. These results provide valuable insights into Rydberg blockade phenomena, reducing the need for precise quantification of Rydberg atoms and offering potential applications in quantum information processing and the investigation of sub-Poissonian atom-number fluctuations in mesoscopic atomic ensembles.

### v2

@@ IV.1 Quantum logic gates
For multiatom ensemble qubits, Saffman’s group demonstrates coherence control between two ensemble qubits and achieves Rydberg blockade between them with a fidelity of 0.89(1).Ebert et al. (2015) However, fluctuations in the collective Rabi frequency could greatly affect the ensemble qubits operation fidelity. Thus, a scheme is proposed to simulate quantum process tomography to characterize Rydberg gates based on collective states using optimal stimulated Raman adiabatic passage pulses. Beterov et al. (2016b) Recently, significant progress has been made in achieving entanglement gates through coherent manipulation of deterministically assembled arrays, employing local optical control to achieve entanglement using interatomic interactions, although individual addressing of atoms remains a challenge experimentally. Zeng et al. (2017); Scholl et al. (2023); Evered et al. (2023); Ma et al. (2023) Additionally, due to the excellent transmission characteristics of photons but the lack of strong interactions, entangling gate schemes for photon-atom hybridization have been proposed. Garcia-Escartin and Chamorro-Posada (2012); Paredes-Barato and Adams (2014); Hao et al. (2015); Das et al. (2016); Sárkány, Fortágh, and Petrosyan (2015) Specifically, the combination of a Rydberg blockade and optical cavities effectively enhances the gate robustness, Das et al. (2016) while a two-photon interference through virtual photon exchanges renders the internal state dynamics of two neutral-atom qubits insensitive to the thermal photon state. Sárkány, Fortágh, and Petrosyan (2015)
@@
Figure 12: Level scheme for EIT storage of the control qubit (a) and target qubit (b). (c) The schematic of the experimental setup. Here, $|R\rangle$ and $|L\rangle$ are used to define polarization states of photons. Reproduced with permission from Tiarks et al., Nature Physics 15(2), 124-126 (2019). Copyright 2019 Springer Nature.Tiarks et al. (2019)
@@
Tiarks et al. achieved the first experimental realization of a photon-photon quantum gate using Rydberg interactions, as documented in their study.Tiarks et al. (2019) By harnessing both EIT and Rydberg interactions within an ultracold atomic ensemble, they effectively encoded quantum information carried by photons into specific atomic states, as depicted in Fig. 12(a). This mapping facilitates the control of photon loss through post-selection, thereby mitigating its detrimental impact on quantum information and ensuring resilience against errors stemming from photon loss. Notably, the photon interaction occurs exclusively when they share the same polarization $|L\rangle$ due to the Rydberg component, and their interaction during temporal and spatial overlap results in the accumulation of a conditional phase shift, thus realizing a controlled phase flip gate, as illustrated in Fig. 12(b). By altering the basis, this gate can be converted into a CNOT gate. The fidelity of the CNOT truth table and the entangling-gate fidelity, post-selected upon the detection of control and target photons, are reported to be 70(8)and 63.7(4.5)%, respectively. This technique employs effectively Rydberg interactions and EIT, allowing for strong, long-range interactions between photons while also solving the difficulty of restricted photon interactions. By mapping Rydberg interactions onto photons, this gate approach stores and manipulates quantum information, demonstrating high fidelity and potential advances in quantum communication systems and quantum networking applications.
@@
Figure 13: Schematic diagram of converting a circuit into an atomic arrangement, which consists of two types of atoms, and superatoms are used as impurities. Reproduced with permission from Cesa et al., Phys. Rev. Lett. 131(17), 170601 (2023). Copyright 2023 American Physical Society.Cesa and Pichler (2023)
@@ IV.2 Quantum entanglement
Quantum entanglement plays a crucial role in quantum computing and quantum information processing. Early experiments of entangled states have focused on photons and ions. Prevedel et al. (2009); Wieczorek et al. (2009); Häffner et al. (2005) However, these schemes face challenges in scaling to multiqubit entanglement and have a low success probability. Rydberg ensembles offer a promising solution to overcome these limitations and realize scalable entanglement of neutral atoms. Atom vapors serve not only as quantum memory, but also as “good” qubits due to strong Rydberg interactions. The theoretical study shows that high-efficiency cluster states can be created via single-qubit measurements with Rydberg ensembles. Zwierz and Kok (2009) Utilizing a fast entanglement generation scheme with a Lipkin-Meshkov-Glick-type Hamiltonian, a 700-atom cat state has been generated with a 50-fold improvement in the size of the entangled state. Khazali (2018) To further enhance fidelity, Gärttner proposes a scheme to herald the entangled state by detecting an ion, achieving near-unity fidelity. Gärttner (2015) Additionally, addressing decoherence concerns, one can combine Zeno dynamics and shortcuts to adiabaticity with Rydberg superatoms, encoding quantum information in the collective state of superatoms to rapidly realize entangled states, including Bell states and $W$ states. Ji et al. (2020) Similarly, fast conversion of three-particle Dicke states to four-particle Dicke states with Rydberg superatoms can be implemented based on Zeno dynamics and shortcuts to adiabaticity, offering high fidelity and robustness against decoherence. Liu et al. (2023b) In Rydberg ensembles, large-scale entangled states can also be prepared theoretically by sending two Rydberg superatoms belonging to two different atomic $W$ states into a cavity without qubit loss. Zhang et al. (2022a)
@@
Figure 14: Entanglement between Rydberg superatom and photon. (a) Schematic diagram of the entanglement of light with optical atomic excitation. Mapping the collective atomic excitation into quantum fields $|\phi\rangle_{A,B}$ by implementing resonance on the $|r\rangle$ and $|e\rangle$ used laser fields $\Omega_{A,B}$. Reproduced with permission from Li et al., Nature 498(7455), 466-469 (2013). Copyright 2013 Springer Nature.Li, Dudin, and Kuzmich (2013) (b) Experimental preparation of semideterministic atom-photon entanglement. Reproduced with permission from Li et al., Phys. Rev. Lett. 123(14), 140504 (2019). Copyright 2019 American Physical Society.Li et al. (2019c)
@@
Photons are excellent carriers for quantum information, and the entanglement between atom and photon holds promise for achieving long-range quantum information distribution. Most ensemble-based atom-photon entanglement schemes are probabilistic and suffer from limited scalability. In 2013, Li et al. introduced a new protocol to generate entanglement between light and optical atomic excitations with intrinsic determinism. Li, Dudin, and Kuzmich (2013) In this scheme, the dephasing of optical atomic coherence is suppressed by state-insensitive confinement of ground and Rydberg states of ultra-cold atomic gas trapped in a one-dimensional optical lattice at 1004 nm. This confinement eliminates the differential energy shift between the ground state and the Rydberg state $\delta U=U_{r}(r)-U_{g}(r)$ and maintains coherence between them. Such a scheme has significant implications for extending the ground-Rydberg coherence and building functional Rydberg quantum networks. Figure 14(a) illustrates the three-step process for the entanglement protocol: (i) Excitation of the atomic ensembles from the ground state $|G\rangle$ to the single excited Rydberg state $|R\rangle$. (ii) Generation of the entanglement state $|R\rangle\rightarrow|R\rangle|0\rangle_{A}+|G\rangle|1\rangle_{A}$ through a retrieval field $\Omega_{A}$. (iii) Verification of the atom-photon entanglement by analyzing the correlations of photoelectric detection events. This scheme is compatible with quantum logic operations and long-term storage of quantum information, paving the way for quantum networks that enable secure communication and distributed quantum computing.
@@
In 2019, an alternative atom-light entanglement protocol was experimentally demonstrated, leveraging Rydberg blockade and Raman coupling techniques. Li et al. (2019c) The manipulation of atoms within a confined volume of approximately 10 $\mu$m in all three dimensions ensures high effectiveness of Rydberg blockade, with an atomic density of approximately $10^{11}/\rm cm^{3}$. In the schematic depicted in Fig. 14(b), the atomic ensemble is initially prepared in a ground state $|g\rangle$, following which Rydberg blockade is employed to sequentially generate two collective excitations—one in a ground state $|s\rangle$ and the other in a Rydberg state $|r\rangle$. These collective excitation states, referred to as two superatoms, possess distinct momenta, are then entangled by incorporating the momentum degree of freedom and interfering them through Raman coupling. Finally, one excitation is retrieved using a read beam, converting the momentum-entangled atomic state into entanglement between the polarization of the single photon and the momentum of the remaining atomic excitation, thereby achieving atom-photon entanglement. The high fidelity measurement of entanglement between the polarization of the single photon and the momentum of the atomic excitation underscores the practicality of employing this approach for quantum information processing tasks like quantum repeaters and networks. Although this scheme is semi-deterministic compared to the approach in Ref. Li, Dudin, and Kuzmich, 2013, it remains immune to phase fluctuations and is devoid of high-order or vacuum components, which are essential for remote entanglement generation. After that, the same group presented a report detailing the deterministic creation of entanglement between an atomic ensemble and a single photon.Sun et al. (2022) This achievement was accomplished through a cyclical process termed “retrieving and patching”, which was applied sequentially to the Rydberg levels. By leveraging the Rydberg blockade effect, they ensured that a full excitation was generated only if a preceding excitation existed in one of the Rydberg levels. This process led to the creation of a photon in a delayed temporal mode, entangled with the Rydberg levels hosting the collective excitation. Verification of the entanglement between the atomic ensemble and the single photon was conducted by retrieving the atomic excitation as a second photon and performing correlation measurements. The results yielded an entanglement fidelity of 87.8%.
@@
The entanglement between photons holds tremendous potential across various fields, such as metrology and quantum computing. A challenge is that interactions between photons are inherently weak. A lot of efforts have been spent on creating sizable photon-photon interactions. Heisenberg and Euler (1936); Busche et al. (2017); Firstenberg et al. (2013) Numerous schemes have been proposed for the preparation of multiphoton entanglement, including spontaneous parametric downconversion Pan et al. (2012); Zhong et al. (2018) and the entangling of single photons through atom-photon interactions. Gheri et al. (1998); Schön et al. (2005) However, these schemes often suffer from drawbacks such as poor scalability of the photon number, complex setups, and other requirements on the system. Recently, Yang et al. proposed a novel scheme to generate multiphoton entanglement with a Rydberg superatom. Yang et al. (2022a) In their approach, an atomic ensemble is used, where an atomic entanglement state $|\Psi\rangle_{\mathrm{a}}$ is prepared by applying collective pulses with the Rydberg blockade mechanism. Subsequently, a “retrieving and patching” sequence is applied to $|r_{1}\rangle$, causing the Rydberg superatom to emit a photon and recreate the excitation. This sequence is repeated for $|r_{2}\rangle$, resulting in the creation of the entanglement state $(|0\rangle_{1}|1\rangle_{2}|L\rangle+|1\rangle_{1}|0\rangle_{2}|E\rangle)/\sqrt{2}$. By repeating these processes $(m-1)$ times and retrieving the two atomic qubits, the GHZ-type multiphoton entanglement $(|L\rangle^{\otimes m}+|E\rangle^{\otimes m})/\sqrt{2}$ can be obtained, where $|L\rangle$ and $|E\rangle$ represent a photon in the late and early modes, respectively. To enhance the performance of multiphoton entanglement, the scheme employs a low-finesse cavity to improve photon collective efficiency and polarizing beamsplitters with a Pockels cell in measurement to minimize exposure. The fidelity of the final prepared three- and six-photon GHZ states is $82.9\pm 0.3\%$ and $61.8\pm 2.6\%$, respectively.
@@
Figure 15: Photonic entanglement filter with Rydberg superatoms. (a) Illustration of Rydberg entanglement filter protocol. (b) Extracting high-fidelity entangled state from input state with low fidelity. Reproduced with permission from Ye et al., Nature Photonics 17(6), 538–543 (2023). Copyright 2023 Springer Nature.Ye et al. (2023)
@@ IV.3 Quantum simulation
Quantum simulation serves as a powerful tool for understanding and predicting many-body dynamics, allowing access to a wide range of phenomena, including thermal equilibrium relaxation in closed systems and non-thermal behaviors in integrable and disordered systems.Daley et al. (2022) A particularly promising platform for investigating the dynamics of strongly interacting many-body quantum systems is provided by coherent laser-driven Rydberg atomic gases.Weimer (2015); Lechner and Zoller (2015); Ding et al. (2024) Currently, spin systems based on Rydberg atoms have emerged as a promising platform for quantum simulation because of their superior interaction flexibility and strength, high controllability, and effective isolation from the environment.
One theoretical strategy suggested by Sun and Robicheaux involves simulating Rydberg gases and exploring two-body correlation by driving the superatoms, which can be subdivided into smaller regions called pseudoatoms, with a laser having a Rabi frequency following a Poissonian distribution. Sun and Robicheaux (2008) This method incorporates the classical correlationZhou et al. (2006) and provides a more accurate depiction of the Rydberg gas compared to the mean field approximation. Using the superatom picture, Olmos and colleagues conducted extensive theoretical research on fermionic collective excitations and thermalization in a symmetric quasi-one-dimensional ring lattice gas of Rydberg atoms. Olmos, González-Férez, and Lesanovsky (2009); Olmos, Müller, and Lesanovsky (2010) In the theoretical studies of Refs. Pohl, Demler, and Lukin, 2010; van Bijnen et al., 2011, the Rydberg superatoms serve as building blocks for the creation of Rydberg crystals using tailored laser excitation schemes. These crystals are used as representations of dilute metallic solids with tunable lattice parameters, allowing access to a wide range of fundamental phenomena in condensed matter physics.
Figure 16: Average density of holes $\rho_{h}$, Mandel $Q$ parameter for the total number of holes, and amplitudes of the density–density correlation function $g^{\left(2\right)}\left(d\right)$ for the periods of $d=2a$ (blue circles) and $d=3a$ (red circles), versus the hopping rate $\Gamma_{t}$. Reproduced with permission from Letscher et al., New J. Phys. 19(11), 113014 (2017). Copyright 2017 Institute of Physics, licensed under a Creative Commons Attribution 3.0 licence.Letscher, Petrosyan, and Fleischhauer (2017)
@@
Figure 17: Ground state population and corresponding Fourier transform in $N$-body ($N=3,4$) configuration. Reproduced with permission from Kim et al., PRX Quantum 1(2), 020323 (2020). Copyright 2020 American Physical Society, licensed under a Creative Commons Attribution 4.0 International license.Kim et al. (2020)
@@
Figure 18: The numerically calculated probabilities $q_{Ry}\left(n\right)$ to excite $n$ Rydberg atoms in the mesoscopic atomic ensemble with $\bar{N}=7$ atoms, randomly distributed in the optical dipole trap with different radius $r$. Reproduced with permission from Beterov et al., Phys. Rev. A 90(4), 043413 (2014). Copyright 2014 American Physical Society.Beterov et al. (2014)
@@
One of the challenges in observing the Rydberg blockade is to obtain a precise measurement of the quantity of Rydberg atoms. Due to low detection efficiencies, accurately determining the number of excited Rydberg atoms becomes problematic, which is essential for exploiting the Rydberg blockade in quantum information applications. Beterov et al. investigated the complex dynamics of Rydberg excitations in atomic ensembles influenced by a classical electromagnetic field.Beterov et al. (2014) They demonstrated the observation of collapses and revivals in the average count of Rydberg excitations, resembling the behavior predicted by the Jaynes-Cummings (JC) model. Through the analysis of randomly loaded optical dipole traps, they predicted the occurrence of collapses and revivals in Rabi oscillations, emphasizing how the oscillation frequency depends on the number of interacting atoms. Moreover, the study examines the impact of finite interaction strengths and laser linewidth on the detectability of the revivals. Considering a randomly loaded optical dipole trap with an average atom number of $\bar{N}=7$, data in Fig. 18 illustrate the excitation dynamics in a Rydberg state in the ensemble. In Fig. 18(a), the simulated dynamics of the probability of single-atom Rydberg excitation $q_{Ry}\left(1\right)$ for $r=2~{}\mu\rm m$ show collapses and revivals, while $q_{Ry}\left(2\right)$ approaches zero, indicating an effective Rydberg blockade. However, with increasing radius of the optical dipole trap, the revivals in $q_{Ry}\left(1\right)$ are diminished, and the probability of exciting two Rydberg atoms $q_{Ry}\left(2\right)$ increases, suggesting the breakdown of the Rydberg blockade. These results provide valuable insights into Rydberg blockade phenomena, reducing the need for precise quantification of Rydberg atoms and offering potential applications in quantum information processing and the investigation of sub-Poissonian atom-number fluctuations in mesoscopic atomic ensembles.
It is noted that significant efforts have been directed towards quantum simulation utilizing Rydberg atom arrays, driven by their exceptional controllability and versatility.Browaeys and Lahaye (2020); Kaufman and Ni (2021) These systems facilitate precise manipulation of individual atoms through laser fields, allowing for the creation of customized potential landscapes and interaction patterns.Omran et al. (2019); Labuhn et al. (2016) Furthermore, the strong and adjustable interactions among Rydberg atoms enable the emulation of various quantum phase transitions.de Léséleuc et al. (2019); Kanungo et al. (2022); Chen et al. (2023a); Bernien et al. (2017); Keesling et al. (2019) The scalability of these arrays, capable of accommodating hundreds of and even moreatoms,Ebadi et al. (2021); Scholl et al. (2021); Bluvstein et al. (2021) offers researchers a robust platform for exploring large quantum systems and addressing computationally challenging problems.Barredo et al. (2016); Endres et al. (2016); Glaetzle et al. (2017); Barredo et al. (2018); Hollerith et al. (2019); Bekenstein et al. (2020); Semeghini et al. (2021); Ebadi et al. (2022); Wu et al. (2022); Bluvstein et al. (2022); Kim et al. (2022); Scholl et al. (2023); Schine et al. (2022); Yan et al. (2022); O’Rourke and Chan (2023); Srakaew et al. (2023); Bornet et al. (2023); Chen et al. (2024) Leveraging unique properties of Rydberg superatoms, we anticipate opportunities for simulating quantum spin models, probing topological phases, and studying thermalization. This can be achieved using techniques such as trapping many atoms per blockade radius in optical lattices or arrays of microscopic dipole traps spaced a few micrometers apart.Côté et al. (2006); Weimer et al. (2010); Viteau et al. (2011); Anderson, Younge, and Raithel (2011); Weber et al. (2015); Gross and Bloch (2017); Whitlock, Glaetzle, and Hannaford (2017); Mei et al. (2022)